# Comparing Bayesian and Frequentist Inference in Biological Models: A Comparative Analysis of Accuracy, Uncertainty, and Identifiability

### v1

@@ Abstract
Mathematical models inform inference and forecasting in ecology and epidemiology, but results depend on the estimation framework. Here, we compare Bayesian and Frequentist approaches across three biological models using four datasets: Lotka–Volterra predator–prey dynamics (Hudson Bay), a generalized logistic model (lung injury and 2022 U.S. mpox), and an SEIUR epidemic model (COVID-19 in Spain). Both approaches use a normal error structure to ensure a fair comparison. We first assessed structural identifiability to determine which parameters can theoretically be recovered from the data. We then evaluated practical identifiability and forecasting performance using four metrics: mean absolute error (MAE), mean squared error (MSE), 95% prediction interval (PI) coverage, and weighted interval score (WIS). For the Lotka–Volterra model with both prey and predator data, we analyzed three observation scenarios: prey only, predator only, and both simultaneously. The Frequentist workflow was implemented using QuantDiffForecast (QDF) in MATLAB, which fits ODE models via nonlinear least squares and quantifies uncertainty through parametric bootstrap. The Bayesian workflow was implemented using BayesianFitForecast (BFF), which employs Hamiltonian Monte Carlo sampling via Stan to generate posterior distributions and diagnostics such as the Gelman–Rubin $\hat{R}$ statistic. Results show that Frequentist inference performs best in well-observed settings with rich data (e.g., GLM for lung injury and mpox, or Lotka–Volterra when both species are observed). In contrast, Bayesian inference excels when latent-state uncertainty is high and data are sparse or partially observed, as seen in the SEIUR model for COVID-19. Structural identifiability analysis clarifies these patterns: full observability enhances both frameworks, while limited observability constrains parameter recovery regardless of the method. This comparative analysis provides practical guidance for selecting inference frameworks based on data richness, observability, and uncertainty quantification needs.
@@ 1 Introduction
Mathematical models based on ordinary differential equations (ODEs) are essential tools for understanding and forecasting dynamics in ecology and epidemiology (modeling_overview; mohammed2024trophiccascade; alsammani2025cholera). Model-based forecasts have been instrumental in managing major public health crises over the past decade. During the COVID-19 pandemic, forecasts guided resource allocation and social distancing policies (dixon2022comparison; cheng2023real; lutz2019applying; shearer2020infectious; chowell2022ensemble; chowell2022sub; alsammani2025cholera). The U.S. CDC’s FluSight Challenge used models to optimize influenza vaccine distribution and public health messaging (reich2019collaborative; mcgowan2019collaborative; biggerstaff2014estimates). During the West African and DRC Ebola outbreaks, models predicted transmission patterns and evaluated intervention effectiveness (chowell2017perspectives; funk2019assessing; meltzer2014estimating; chretien2015mathematical; chowell2014west; roosa2020multi). More recently, forecasting models for mpox predicted its spread and assessed containment measures (bleichrodt2023real; bleichrodt2023retrospective; charniga2024nowcasting; chowell2024growthpredict). In ecology, predator-prey models help predict population dynamics and inform conservation strategies (lotka_volterra_data). These applications demonstrate that reliable forecasts depend on accurate parameter estimation (forecast_reliability; Chowell2024QuantDiffForecast; Karami2025BayesianFitForecast).
@@
However, the ability to recover meaningful parameters from data critically depends on identifiability—the question of whether model parameters can be uniquely determined from available observations (identifiability_review; practical_identifiability; atanasov2025honeybee). Identifiability can fail when data are sparse, models are over-parameterized, or parameters are strongly correlated, leading to misleading inferences and unreliable forecasts. Recent work has shown that factors such as vaccination behavior, immunity duration, and data completeness shape both epidemic trajectories and the reliability of parameter estimates (Alsammani2025MBE). Understanding how identifiability interacts with estimation frameworks is essential to ensure that model-based forecasts are interpretable and reproducible. Structural or practical identifiability limitations can cause two inference frameworks to produce different results even under identical models and data.
@@
Two estimation paradigms dominate the field: Bayesian and Frequentist methods. Frequentist methods typically calibrate ODE models by optimizing a likelihood function or minimizing an objective function such as the sum of squared differences between observed and predicted values (gneiting2008probabilistic; mwambi2011frequentist; chowell2017fitting; chowell2020real). These methods use algorithms like gradient descent or the Levenberg-Marquardt algorithm, assume specific distributions for measurement errors (e.g., Gaussian or Poisson), and quantify uncertainty through bootstrapping techniques (banks2014modeling; pruitt2024role; transtrum2012optimal; huang2024nonlinear; bates1988nonlinear; cao2012penalized; ramsay2017dynamic; seber2003nonlinear). Frequentist methods are computationally efficient and often perform well when data are abundant and high quality (deterministic_text; Chowell2024QuantDiffForecast). The QuantDiffForecast (QDF) toolbox implements this workflow for fitting models and generating predictions with quantified uncertainty (chowell2024parameter).
@@
Bayesian methods, in contrast, apply Bayes’ theorem to combine prior distributions of parameters with the likelihood of observed data, producing posterior distributions that explicitly incorporate uncertainty (greenland2009bayesian; mckinley2014simulation; kypraios2017tutorial; girolami2008bayesian; grinsztajn2021bayesian; bouman2024bayesian; gelman2020bayesian; belasso2023bayesian). These methods typically use Markov Chain Monte Carlo (MCMC) algorithms to approximate posterior distributions, providing comprehensive measures of parameter uncertainty and credible intervals (martin2020computing; dunson2001commentary; harel2018multiple; grinsztajn2006bayesian; annis2017bayesian; kelter2020bayesian). Bayesian methods can better navigate complex parameter spaces, avoid local optima, and handle incomplete or noisy data (gelman1996physiological; huang2006hierarchical; huang2020bayesian). Tools like Stan facilitate Bayesian estimation and forecasting, allowing rigorous uncertainty quantification and model validation (bayesian_text; gelman_rhat; posterior_diagnostics; Karami2025BayesianFitForecast; monnahan2017faster; burkner2017brms). The BayesianFitForecast (BFF) toolbox implements this workflow with diagnostics such as the Gelman-Rubin $\hat{R}$ statistic.
@@
Despite extensive use of both paradigms, prior comparisons often vary models, likelihoods, or preprocessing between methods, making it difficult to attribute observed differences to the estimation framework itself (benchmarking_pitfalls). To address this gap, we conduct a controlled comparison of Bayesian and Frequentist inference under standardized conditions: the same models, the same normal error structure, and harmonized data preprocessing. We analyze three systems and four datasets representing different levels of complexity and observability: the Lotka-Volterra (LV) predator-prey model (Hudson Bay lynx-hare data), a generalized logistic model (GLM) for lung injury and the 2022 U.S. mpox outbreak, and an SEIUR epidemic model for the first COVID-19 wave in Spain (lotka_volterra_data; lung_injury_glm; mpox_us_2022; spain_covid_firstwave). For the LV model, we analyze three observation scenarios (prey only, predator only, and both simultaneously) to assess how partial observability affects parameter recovery.
@@
We integrate structural identifiability analysis to determine which parameters can theoretically be recovered from the data, separating fundamental data limits from algorithmic limits (identifiability_methods; stigter2023structural; dankwa2022identifiability). We then evaluate practical identifiability and forecasting performance using four metrics: mean absolute error (MAE), mean squared error (MSE), 95% prediction interval (PI) coverage, and weighted interval score (WIS) (wis_ref; coverage_ref; mcmc_convergence; Chowell2024QuantDiffForecast; Karami2025BayesianFitForecast; Gneiting2007).
@@
Our objectives are to: (i) compare prediction accuracy across Bayesian and Frequentist inference, (ii) assess uncertainty calibration and diagnostics, and (iii) relate observed performance to structural versus practical identifiability under full versus partial observation. This provides practical guidance on when each paradigm is preferable based on data richness, observability, and uncertainty quantification needs (Karami2025BayesianFitForecast; Chowell2024QuantDiffForecast).
@@
The remainder of this paper is organized as follows. Section 2 describes the four datasets analyzed in this study, including the Hudson Bay lynx-hare data, lung injury outbreak data, mpox epidemic data, and COVID-19 data from Spain. Section 3 presents the three mathematical models employed: the Lotka–Volterra predator-prey model, the generalized logistic model, and the SEIUR epidemic model. Section 4 details the Bayesian and Frequentist inference methodologies, including likelihood specifications, prior distributions, uncertainty quantification procedures, and performance metrics. Section 5 presents the structural identifiability analysis for all models under different observation scenarios, establishing theoretical bounds on parameter recovery. Section 6 reports empirical results, including parameter estimates, performance metrics, and convergence diagnostics for all models and datasets. In Section 7, the Discussion synthesizes our findings, interprets the role of structural identifiability in explaining performance differences, and provides practical guidance for method selection. The Conclusion, Section 8, summarizes the main contributions and implications for biological modeling practice.
@@ 2.1 Hudson Bay Lynx–Hare Data
The Hudson Bay lynx–hare dataset consists of annual counts of Canadian lynx (Lynx canadensis) and snowshoe hare (Lepus americanus) populations from 1900 to 1920, derived from Hudson’s Bay Company pelt return records (carpenter2018predator; elton1924periodic). These records serve as a proxy for population abundance, with pelt counts reflecting relative population size. The dataset contains $N=21$ yearly observations for both prey (hare) and predator (lynx) populations. This dataset is particularly valuable for assessing parameter identifiability in predator–prey dynamics because it provides simultaneous observations of both interacting species over multiple population cycles. The data have been widely used as a benchmark for testing ecological models and parameter estimation methods. No smoothing, interpolation, or imputation was applied to the original data.
@@ 2.2 Lung Injury Data (EVALI)
This dataset comprises weekly incident case counts of electronic-cigarette or vaping product use-associated lung injury (EVALI) reported in the United States during 2019 (krishnasamy2020evali; lozier2019evaliupdate; chowell2020evalimodel). The outbreak was first identified in mid-June 2019, with cases rapidly escalating through the summer and fall before declining by November 2019. We analyzed $N\approx 21$ weeks of data spanning mid-June to early November 2019, as reported through the Centers for Disease Control and Prevention (CDC) Morbidity and Mortality Weekly Report (MMWR) surveillance system. Case definitions followed CDC criteria for confirmed and probable EVALI cases. The data were aggregated by epidemiological week (MMWR week) with no backfill corrections applied beyond those included in the official CDC reports. This dataset represents a short-duration outbreak with a clear peak, making it suitable for testing phenomenological growth models under well-defined epidemic dynamics.
@@ 2.3 Mpox Data (United States, 2022)
We analyzed weekly incident confirmed or probable mpox cases in the United States during the 2022–2023 outbreak as reported in (cdc2022mpox; chiu2023mpox). The outbreak began in May 2022 and peaked in August 2022 before declining through early 2023. Data were compiled from CDC national surveillance reports aggregated by MMWR week. Case definitions followed CDC guidelines for confirmed (laboratory-confirmed orthopoxvirus with specimen typed as monkeypox virus) and probable cases (epidemiologically linked to a confirmed case without laboratory confirmation). The temporal window analyzed corresponds to the public reporting period documented in CDC surveillance dashboards (see references for access dates). This dataset provides an example of an emerging infectious disease outbreak with intensive surveillance and public health response, allowing evaluation of model performance during rapidly evolving epidemic conditions.
@@ 2.4 COVID-19 Data (Spain, First Wave)
This dataset consists of daily incident laboratory-confirmed COVID-19 cases reported during the first epidemic wave in Spain from February through May 2020 (espana2020covid; chowell2020covidspain). Data were compiled from official reports by the Spanish Ministry of Health (Ministerio de Sanidad), with case dates based on report date rather than symptom onset date due to data availability constraints during the early pandemic response. The first wave in Spain exhibited rapid exponential growth followed by a peak in late March 2020 and subsequent decline following implementation of strict non-pharmaceutical interventions including a national lockdown. We restricted analysis to the first wave period to avoid complications from changing surveillance protocols, intervention policies, and emergence of new variants that occurred in later waves. Known reporting artifacts explicitly flagged in official data releases were excluded. This dataset represents a large-scale epidemic with significant underreporting and latent compartments (exposed and unreported infectious individuals), making it ideal for assessing inference methods under partial observability and model complexity.
@@ 2.5 Preprocessing Summary
2.5.0.1 Temporal aggregation
The Hudson Bay lynx–hare data were used as published without aggregation. Lung injury and mpox data were aggregated at the weekly level (MMWR epidemiological week) as reported in official CDC surveillance. COVID-19 data for Spain were analyzed at the daily level as reported by the Spanish Ministry of Health.
2.5.0.2 Data type
The Hudson Bay dataset represents population abundance (pelt counts), while all epidemic datasets (lung injury, mpox, COVID-19) represent incident case counts per reporting period. For the generalized logistic model (GLM), cumulative case counts were derived by summing incident counts, and model predictions were compared to incident data via the time derivative of cumulative cases.
2.5.0.3 Missing values
No missing values were present in the final temporal windows used for model fitting. No imputation was performed.
2.5.0.4 Smoothing
No smoothing was applied to the data used for parameter estimation or forecasting. Figures may display smoothed overlays for visualization purposes only; these smoothed curves were not used in model fitting.
2.5.0.5 Temporal windows
Modeling windows for calibration and forecasting were selected based on epidemic phase and data availability. Specific windows for each dataset are described in the Results section.
2.5.0.6 Data quality control
For COVID-19 data, known reporting artifacts explicitly flagged in official Spanish Ministry of Health releases were excluded. For all other datasets, data were used as published in official surveillance reports without additional corrections.
Figure 1: Time series of ecological and epidemic population dynamics. The Hudson Bay Lynx–Hare dataset shows annual prey and predator abundance (1900–1920), while the remaining panels depict weekly or daily reported cases from major disease outbreaks: Lung Injury (Electronic cigarette or vaping product use-associated lung injury (EVALI), US, 2019), Mpox (US, 2022), and COVID-19 (Spain, 2020).
@@ 3 Models
In this study, we employed three compartmental biological and epidemiological models to comprehensively assess the parameter identifiability of our two approaches: (i) the LV model, describing predator-prey interactions and population feedback mechanisms; (ii) the Generalized Logistic Model, capturing flexible epidemic growth through nonlinear case dynamics; and (iii) the SEIUR model, explicitly tracking susceptible, exposed, infectious (reported and unreported), and recovered populations to account for both observed and hidden transmission. These models were selected to represent a progression in model complexity and data observability: from a fully observed ecological system (LV), to a parsimonious single-equation epidemic model (GLM), to a multi-compartment latent-state model with partial observability (SEIUR). This progression enables a systematic comparison of Bayesian and Frequentist inference performance across increasing levels of model structure and latent uncertainty.
3.0.0.1 LV model
The LV equations (Lotka 1925; Volterra 1926, 1927) consist of a pair of first-order ODE describing the population dynamics of two interacting species: one predator and one prey. Let $x$ denote the prey population and $y$ the predator population at time $t$. Volterra modeled the temporal dynamics of these populations as follows:

### v2

@@ Abstract
Mathematical models are widely used to study ecological and epidemiological systems, but the accuracy of resulting inferences and forecasts depends on the estimation framework. This study compares Bayesian and Frequentist approaches across three biological models using four datasets: the Lotka–Volterra (LV) predator–prey system, a generalized logistic model (GLM) applied to lung injury and mpox data, and SEIUR epidemic model for COVID-19. To ensure a fair comparison, both approaches were implemented with a normal error structure. We first examined structural identifiability to determine which parameters can be recovered in principle. We then evaluated practical identifiability and fitting performance using four metrics: mean absolute error (MAE), mean squared error (MSE), prediction interval (PI) coverage, and weighted interval score (WIS). For the Lotka–Volterra model, we studied three types of observations: prey only, predator only, and both species together. The Frequentist workflow, implemented through QuantDiffForecast, uses nonlinear least squares and parametric bootstrap to quantify uncertainty. The Bayesian workflow, implemented through BayesianFitForecast, uses Hamiltonian Monte Carlo sampling to obtain posterior distributions and diagnostic measures. Our results show that Frequentist inference performs well when the system is fully observed data, as in the GLM or in the Lotka–Volterra model when both species are observed. Bayesian inference performs better when uncertainty is high and observations are limited or indirect, as seen in the SEIUR epidemic model. The identifiability analysis helps explain these differences by showing how observability shapes the reliability of parameter recovery and uncertainty quantification.
@@ 1 Introduction
Mathematical models based on ordinary differential equations (ODEs) are essential tools for understanding and forecasting dynamics in ecology and epidemiology [1, 2, 3]. Model-based forecasts have been instrumental in managing major public health crises over the past decade. During the COVID-19 pandemic, forecasts guided resource allocation and social distancing policies [4, 5, 6, 7, 8, 9, 91].
The U.S. CDC’s FluSight Challenge used models to optimize influenza vaccine distribution and public health messaging [10, 11, 12]. During the West African and Democratic Republic of the Congo (DRC) Ebola outbreaks, models predicted transmission patterns and evaluated intervention effectiveness [13, 14, 15, 16, 17, 18]. More recently, forecasting models for mpox predicted its spread and assessed containment measures [19, 20, 21, 22].
In ecology, predator–prey models help predict population dynamics and inform conservation strategies [23]. These applications demonstrate that reliable forecasts depend on accurate parameter estimation [24, 25, 26]. However, the ability to recover meaningful parameters from data critically depends on identifiability, that is, whether model parameters can be uniquely determined from available observations [27, 28, 29].
@@
Identifiability can fail when data are sparse, models are over-parameterized, or parameters are strongly correlated, leading to misleading inferences and unreliable forecasts. Recent work has shown that factors such as vaccination behavior, immunity duration, and data completeness shape both epidemic trajectories and the reliability of parameter estimates [30]. Understanding how identifiability interacts with estimation frameworks is essential to ensure that model-based forecasts are interpretable and reproducible. Structural or practical identifiability limitations can cause two inference frameworks to produce different results even under identical models and data.
@@
Two estimation paradigms dominate the field: Bayesian and Frequentist methods. Frequentist methods typically calibrate ODE models by optimizing a likelihood function or minimizing an objective function such as the sum of squared differences between observed and predicted values [31, 32, 33, 34]. These methods use algorithms such as gradient descent or the Levenberg–Marquardt algorithm, assume specific distributions for measurement errors (e.g., Gaussian, Poisson, or Normal), and quantify uncertainty through bootstrapping techniques [35, 36, 37, 38, 39, 40, 41, 42]. Frequentist methods are computationally efficient and often perform well when data are abundant and high quality [43, 25]. The QuantDiffForecast (QDF) toolbox implements this workflow for fitting models and generating predictions with quantified uncertainty [25].
@@
Bayesian methods, in contrast, apply Bayes’ theorem to combine prior distributions of parameters with the likelihood of observed data, producing posterior distributions that explicitly incorporate uncertainty [44, 45, 46, 47, 48, 49, 50, 51]. These methods typically use Markov Chain Monte Carlo (MCMC) algorithms to approximate posterior distributions, providing comprehensive measures of parameter uncertainty and prediction intervals [52, 53, 54, 55, 56, 57]. Bayesian methods can better navigate complex parameter spaces, avoid local optima, and handle incomplete or noisy data [58, 59, 60]. Tools such as Stan facilitate Bayesian estimation and forecasting, allowing rigorous uncertainty quantification and model validation [61, 62, 63, 64, 65].
@@
The BayesianFitForecast (BFF) toolbox implements this workflow with diagnostics such as the Gelman–Rubin $\hat{R}$ statistic [26]. Despite extensive use of both paradigms, prior comparisons often vary models, likelihoods, or preprocessing between methods, making it difficult to attribute observed differences to the estimation framework itself [66]. To address this gap, we conduct a controlled comparison of Bayesian and Frequentist inference under standardized conditions: the same models, the same normal error structure, and harmonized data preprocessing.
We analyze three systems and four datasets representing different levels of complexity and observability: the Lotka–Volterra (LV) predator–prey model (Hudson Bay lynx–hare data), a generalized logistic model (GLM) for lung injury and the 2022 United States mpox outbreak, and an SEIUR epidemic model for the first COVID-19 wave in Spain [67, 68, 69]. For the Lotka–Volterra model, we analyze three observation scenarios (prey only, predator only, and both simultaneously) to assess how partial observability affects parameter recovery.
@@
We integrate structural identifiability analysis to determine which parameters can theoretically be recovered from the data, separating fundamental data limitations from algorithmic limitations [70, 71, 72]. We then evaluate practical identifiability and forecasting performance using four metrics: mean absolute error (MAE), mean squared error (MSE), 95% prediction interval (PI) coverage, and weighted interval score (WIS) [73, 74, 75, 76, 77].
@@
Our objectives are to (i) compare prediction accuracy across Bayesian and Frequentist inference, (ii) assess uncertainty calibration and diagnostics, and (iii) relate observed performance to structural versus practical identifiability under full versus partial observation. This provides practical guidance on when each paradigm is preferable based on data richness, observability, and uncertainty quantification needs [26].
@@
The remainder of this paper is organized as follows. Section 2 describes the four datasets analyzed in this study. Section 3 presents the mathematical models employed. Section 4 details the Bayesian and Frequentist inference methodologies. Section 5 presents the structural identifiability analysis. Section 6 reports the empirical results. The Discussion synthesizes our findings, and the Conclusion summarizes the main contributions and implications for biological modeling practice.
@@ 2.1 Hudson Bay Lynx–Hare Data
The Hudson Bay lynx–hare dataset consists of annual counts of Canadian lynx (Lynx canadensis) and snowshoe hare (Lepus americanus) populations from 1900 to 1920, derived from Hudson’s Bay Company pelt return records [76, 77] . These records serve as a proxy for population abundance, with pelt counts reflecting relative population size. The dataset contains $N=21$ yearly observations for both prey (hare) and predator (lynx) populations. This dataset is particularly valuable for assessing parameter identifiability in predator–prey dynamics because it provides simultaneous observations of both interacting species over multiple population cycles. The data have been widely used as a benchmark for testing ecological models and parameter estimation methods. No smoothing, interpolation, or imputation was applied to the original data.
@@ 2.2 Lung Injury Data (EVALI)
This dataset comprises weekly incident case counts of electronic-cigarette or vaping product use-associated lung injury (EVALI) reported in the United States during 2019 [78, 79, 80] . The outbreak was first identified in mid-June 2019, with cases rapidly escalating through the summer and fall before declining by November 2019. We analyzed $N\approx 21$ weeks of data spanning mid-June to early November 2019, as reported through the Centers for Disease Control and Prevention (CDC) Morbidity and Mortality Weekly Report (MMWR) surveillance system. Case definitions followed CDC criteria for confirmed and probable EVALI cases. The data were aggregated by epidemiological week (MMWR week) with no backfill corrections applied beyond those included in the official CDC reports. This dataset represents a short-duration outbreak with a clear peak, making it suitable for testing phenomenological growth models under well-defined epidemic dynamics.
@@ 2.3 Mpox Data (United States, 2022)
We analyzed weekly incident confirmed or probable mpox cases in the United States during the 2022–2023 outbreak as reported in [81, 82]. The outbreak began in May 2022 and peaked in August 2022 before declining through early 2023. Data were compiled from CDC national surveillance reports aggregated by MMWR week. Case definitions followed CDC guidelines for confirmed (laboratory-confirmed orthopoxvirus with specimen typed as monkeypox virus) and probable cases (epidemiologically linked to a confirmed case without laboratory confirmation). The temporal window analyzed corresponds to the public reporting period documented in CDC surveillance dashboards (see references for access dates). This dataset provides an example of an emerging infectious disease outbreak with intensive surveillance and public health response, allowing evaluation of model performance during rapidly evolving epidemic conditions.
@@ 2.4 COVID-19 Data (Spain, First Wave)
This dataset consists of daily incident laboratory-confirmed COVID-19 cases reported during the first epidemic wave in Spain from February through May 2020 [83, 84] . Data were compiled from official reports by the Spanish Ministry of Health (Ministerio de Sanidad), with case dates based on report date rather than symptom onset date due to data availability constraints during the early pandemic response. The first wave in Spain exhibited rapid exponential growth followed by a peak in late March 2020 and subsequent decline following implementation of strict non-pharmaceutical interventions including a national lockdown. We restricted analysis to the first wave period to avoid complications from changing surveillance protocols, intervention policies, and emergence of new variants that occurred in later waves. Known reporting artifacts explicitly flagged in official data releases were excluded. This dataset represents a large-scale epidemic with significant underreporting and latent compartments (exposed and unreported infectious individuals), making it ideal for assessing inference methods under partial observability and model complexity.
@@ 2.5 Preprocessing summary:
  1. 1.
    Temporal aggregation: The Hudson Bay lynx–hare data were used as published without aggregation. Lung injury and mpox data were aggregated at the weekly level (MMWR epidemiological week) as reported in official CDC surveillance. COVID-19 data for Spain were analyzed at the daily level as reported by the Spanish Ministry of Health.
  2. 2.
    Data type: The Hudson Bay dataset represents population abundance (pelt counts), while all epidemic datasets (lung injury, mpox, COVID-19) represent incident case counts per reporting period. For the generalized logistic model (GLM), cumulative case counts were derived by summing incident counts, and model predictions were compared to incident data via the time derivative of cumulative cases.
  3. 3.
    Missing values: No missing values were present in the final temporal windows used for model fitting. No imputation was performed.
  4. 4.
    Smoothing: No smoothing was applied to the data used for parameter estimation or forecasting. Figures may display smoothed overlays for visualization purposes only; these smoothed curves were not used in model fitting.
  5. 5.
    Temporal windows: Modeling windows for calibration and forecasting were selected based on epidemic phase and data availability. Specific windows for each dataset are described in the Results section.
  6. 6.
    Data quality control: For COVID-19 data, known reporting artifacts explicitly flagged in official Spanish Ministry of Health releases were excluded. For all other datasets, data were used as published in official surveillance reports without additional corrections.
Figure 1: Time series of ecological and epidemic population dynamics. The Hudson Bay Lynx–Hare dataset shows annual prey and predator abundance (1900–1920), while the remaining panels depict weekly or daily reported cases from major disease outbreaks: Lung Injury (EVALI, US, 2019), Mpox (US, 2022), and COVID-19 (Spain, 2020).
@@ 3 Models
In this study, we employed three compartmental biological and epidemiological models to comprehensively assess the parameter identifiability of our two approaches: (i) the Lotka–Volterra model, describing predator-prey interactions and population feedback mechanisms; (ii) the Generalized Logistic Model, capturing flexible epidemic growth through nonlinear case dynamics; and (iii) the SEIUR model, explicitly tracking susceptible, exposed, infectious (reported and unreported), and recovered populations to account for both observed and hidden transmission. These models were selected to represent a progression in model complexity and data observability: from a fully observed ecological system (LV), to a parsimonious single-equation epidemic model (GLM), to a multi-compartment latent-state model with partial observability (SEIUR). This progression enables a systematic comparison of Bayesian and Frequentist inference performance across increasing levels of model structure and latent uncertainty.
3.0.0.1 Lotka-Volterra model
The Lotka–Volterra equations (Lotka 1925; Volterra 1926, 1927) consist of ODEs describing the population dynamics of two interacting species: one predator and one prey. Let $x$ denote the prey population and $y$ the predator population at time $t$. Volterra modeled the temporal dynamics of these populations as follows: